OpenStudy (anonymous):

Angle A and angle B are suplementary angles. Find angle a and angle b. Are my answers right? angle a= (x+50) angle b= (x+100) My answers- angle a=130 degrees angle b= 80 degrees

OpenStudy (anonymous):

no has to add up to 180

OpenStudy (anonymous):

Ugh.. So are you supposed to set them each to 180 degrees?

OpenStudy (anonymous):

x equals 15

OpenStudy (anonymous):

15+50 = 65 15+100=115 115+65= 180